# Exploring Large Language Models for Feature Selection: A Data-centric Perspective

### v1

@@ 6 Outlook
Synergy of LLMs-based and traditional feature selection. As we discuss in Section 1 and 4.2, text-based feature selection with LLMs is competitive and resource-efficient compared with traditional feature selection methods. However, each approach relies on different sources of information—specific samples or context descriptions—to perform feature selection. This diversity in information utilization makes them complementary. It would be valuable to explore how to combine text-based and traditional feature selection methods to create more effective and robust feature selection systems across various data availability scenarios.
Data-driven analysis with Agentic LLMs. In Section 4.2, we conclude that poor statistical inference capabilities in long-sequence input hinder LLMs in data-driven feature selection. While this finding implies the sole adaptation of LLMs may not be enough for performing data-driven feature selection, the introduction of agent-based LLMs should be considered as an alternative (Xi et al., 2023; Wang et al., 2024a). These methods equip LLM with various tools (Paranjape et al., 2023; Yang et al., 2024; Schick et al., 2024) and APIs (Patil et al., 2023; Liu et al., 2024c), enabling them to execute actions and plans to solve complex and multi-step problems. However, there are only a few works that focus on the development of agentic LLMs as data engineers and analytics (Hong et al., 2024; Fang et al., 2024), for actively performing various features or data processing with the assistance of statistical tools or software. Research in this direction will be valuable for enhancing and evaluating LLMs from analytical and statistical perspectives.
Foundation models for feature/data engineering. Many recent works have developed various foundation models in many data mining and machine learning fields, such as graph learning (Liu et al., 2023; Mao et al., 2024; Xia et al., 2024) and time series prediction (Rasul et al., 2023; Jin et al., 2023). A large foundation model for feature/ data engineering should be able to understand different types of information from the datasets and perform efficient manipulation and processing (Cui et al., 2024) to prepare appropriate data for downstream models/ applications. Developing such a foundation model would greatly benefit the data mining and machine learning communities by providing a unified, easy-to-use interface for complex data processing tasks.
@@ 7 Conclusion
In this study, we explore feature selection methods based on LLMs from a data-centric perspective. We categorize existing LLM-based feature selection approaches into two main types: data-driven, which relies on statistical inference from specific samples, and text-based, which utilizes the extensive knowledge of LLMs for semantic association. Our experiments and analyses reveal that text-based feature selection with LLMs outperforms data-driven methods in terms of effectiveness, stability, and robustness. Based on these findings, we introduce a Retrieval-Augmented Feature Selection (RAFS) method designed to manage large volumes of domain-specific feature candidates in the context of cancer survival time prediction. Additionally, we provide a comprehensive analysis of the current challenges and potential opportunities at the intersection of LLMs and feature selection/engineering, aiming to offer deeper insights and guidance for future research in this area.
@@ Appendix A Detailed Instruction
You can have as much text here as you want. The main body must be at most $8$ pages long. For the final version, one more page can be added. If you want, you can use an appendix like this one.
The \$\mathtt{\backslash onecolumn}$ command above can be kept in place if you prefer a one-column appendix, or can be removed if you prefer a two-column appendix. Apart from this possible change, the style (font size, spacing, margins, page numbering, etc.) should be kept the same as the main body.

### v2

@@ 6 Outlook
Synergy of LLMs-based and traditional feature selection. As we discuss in Sections 1 and 4.2, text-based feature selection with LLMs is competitive and resource-efficient compared with traditional feature selection methods. However, each approach relies on different sources of information—specific samples or context descriptions to perform feature selection. This diversity in information utilization makes them complementary. It would be valuable to explore how to combine text-based and traditional feature selection methods to create more effective and robust feature selection systems across various data availability scenarios. Also, it would be interesting to explore the synergy of text-based and data-driven methods to further enhance LLMs-based feature selection under resource constrains.
Data-driven analysis with Agentic LLMs. In Section 4.2, we conclude that poor statistical inference capabilities in long-sequence input hinder LLMs in data-driven feature selection. While this finding implies the sole adaptation of LLMs may not be enough for performing data-driven feature selection, the introduction of agent-based LLMs should be considered as an alternative [68, 61]. These methods equip LLM with various tools [47, 70, 51] and APIs [48, 41], enabling them to execute actions and plans to solve complex and multi-step problems. However, there are only a few works that focus on the development of agentic LLMs as data engineers and analytics [25, 14, 59], for actively performing various features or data processing with the assistance of statistical tools or software. Research in this direction will be valuable for enhancing and evaluating LLMs from analytical and statistical perspectives.
Foundation models for feature/data engineering. Many recent works have developed various foundation models in many data mining and machine learning fields, such as graph learning [38, 43, 69] and time series prediction [49, 27]. A large foundation model for feature/ data engineering should be able to understand different types of information from the datasets and perform efficient manipulation and processing [9] to prepare appropriate data for downstream models/ applications. Developing such a foundation model would greatly benefit the data mining and machine learning communities by providing a unified, easy-to-use interface for complex data processing tasks.
@@ 7 Conclusion
In this study, we explore feature selection methods based on LLMs from a data-centric perspective. We categorize existing LLM-based feature selection approaches into two main types: data-driven, which relies on statistical inference from specific samples, and text-based, which utilizes the extensive knowledge of LLMs for semantic association. Our experiments and analyses reveal that text-based feature selection with LLMs outperforms data-driven methods in terms of effectiveness, stability, and robustness. Based on these findings, we introduce a Retrieval-Augmented Feature Selection (RAFS) method designed to manage large volumes of domain-specific feature candidates in the context of cancer survival time prediction. Additionally, we provide a comprehensive analysis of the current challenges and potential opportunities at the intersection of LLMs and feature selection/engineering in Section 6, aiming to offer insights and guidance for future research in this area.
Acknowledgments
The material in this presentation is based upon work supported by the U.S. Department of Homeland Security under Grant Award Number, 17STQAC00001-08-00, the U.S. Office of Naval Research (ONR) under grant N00014-21-1-4002, and the National Science Foundation (NSF) under grants IIS-2229461. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security and the National Science Foundation.